### v1

@@ 1 Introduction
This work is concerned with the study of resonating behavior arising in wave-scattering problems, which are modeled in the frequency domain through the Helmholtz equation. It is well known from Fredholm theory that, as long as the proper radiation conditions are enforced for closure far away from the scatterer(s), scattering problems admit a unique solution at any nonzero real frequency. However, the well-posedness of the problem is sometimes at risk as a consequence of what we informally call “almost resonances”, namely, frequencies where the inf-sup constant of the problem becomes dangerously close to 0.
@@
Almost-resonance problems are usually exacerbated by large frequencies, since, in the large-frequency regime, the complex resonances of the scattering problem get arbitrarily close to the real axis. Thus, the behavior of the scattered wave can truly get close to resonance. A good understanding of these phenomena is paramount in many fields, e.g., when resonance is sought as a way to increase the power of a signal or, on the opposite end, when one tries to avoid resonance in the context of making a physical system more stable or well-behaved. This justifies our interest in this problem.
@@
Incidentally, we also mention that almost resonances have important consequences on the stability of the scattering problem, e.g., when solved in approximated form by a numerical technique. As such, spotting almost resonances may be of use in understanding the behavior of numerical methods, e.g., in terms of convergence of iterative linear-system solvers.
@@
At almost-resonating frequencies, the scattered wave may be rather large (in some norm) even for small incident waves. This issue may be equivalently stated in terms of a large operator norm (in suitable spaces) of the solution operator of the scattering problem, which, at fixed frequency, maps incident wave to scattered wave. In this work, we use this norm, under the label of “field amplification”, as a quantitative measure of almost-resonating behavior.
@@
For a given scattering problem, our ultimate task is evaluating the field amplification over a prescribed range of frequency values. A key observation is that the cost of evaluating field amplification over a whole frequency range may be rather high, especially if both spatial- and frequency-domain discretizations are fine. Indeed, one is required to assemble the discrete solution operator and then find its norm (e.g., by SVD) at each frequency belonging to a grid that spans the target frequency range.
@@
To lessen this burden, we rely on surrogate modeling, which suggests to replace the expensive high-fidelity scattering model with a more parsimonious low(er)-fidelity surrogate. The construction of such surrogate (the so-called “training” phase) is generally not cheap, since it must involve several evaluations of the high-fidelity model. However, it must be carried out only once, independently of the desired number of frequency-domain samples. Ignoring further training costs for simplicity, the complexity of a surrogate-based method for approximating scattering problems is often $\mathcal{O}\left(c_{\textup{high-fidelity}}n+c_{\textup{surrogate}}N\right)$, with $n$ denoting the numbers of training samples and $N$ denoting the number of target frequency evaluations. Compared to the “brute force” approach where the high-fidelity model is sequentially evaluated, computational time can be saved as long as the sampling is parsimonious ($n\ll N$) and the surrogate is efficient ($c_{\textup{surrogate}}\ll c_{\textup{high-fidelity}}$).
@@
In recent years, surrogate modeling has often been applied with the objective of reducing the simulation costs linked to scattering problems [20, 32, 36, 6, 27, 16, 12, 7]. In order to build an effective surrogate, a careful balance usually has to be struck between a priori knowledge of the high-fidelity problem’s structure and a posteriori information obtained from data. Both a priori and a posteriori information must be leveraged in the quest for an efficient method, i.e., keeping $n$ and $c_{\textup{surrogate}}$ low while achieving an acceptable approximation accuracy. The a priori knowledge on the problem is crucial for this, since it is what enables the selection of the “approximation class” within which the surrogate is then sought through a data-driven strategy. In the context of surrogate modeling, the approximation class often “makes or breaks” the effectiveness of a numerical algorithm.
@@
However, the effectiveness of the algorithm in [12] is ultimately limited by the fact that the chosen approximation class only partially exploits the structure of the high-fidelity problem. Indeed, the method exploits the piecewise-smooth nature of the field amplification but not the meromorphy of the underlying scattering problem, cf. Section 3. In this work, we try to remedy this, by developing methods that better profit from both features. To this aim, we rely on some valuable results from operator theory to draw a key connection between our approximation target, namely, field amplification, and a certain nonlinear eigenvalue problem induced by the scattering problem’s differential operator.
A key aspect of our work is that, in building a surrogate for the field amplification, we assume that only purely real wavenumbers can be considered and sampled. Fundamentally, we do this because field amplification is a real quantity, related to real wavenumbers, so it should be possible to approximate it by sampling real wavenumbers only. However, and from a more practical viewpoint, non-purely-real wavenumbers are sometimes out of reach due to limitations in the numerical software that is used to discretize the scattering problem in the first place, and ultimately to compute field amplification. Indeed, adding an imaginary component to the wavenumber is arguably nonphysical, so that, by design, it may not be allowed by, e.g., finite- or boundary-elements libraries.
@@ 2 Boundary-integral equations
The unknown is the total wave $u_{\textup{tot}}:\mathbb{R}^{d}\to\mathbb{C}$, $\textup{i}=\sqrt{-1}$ is the imaginary unit, and $u_{\textup{inc}}:\mathbb{R}^{d}\to\mathbb{C}$ is a given incident wave. Moreover, $n:\mathbb{R}^{d}\to\mathbb{R}_{>0}$ is the index of refraction, which is assumed to attain a constant value inside and outside of a suitable bounded, Lipschitz domain $\Omega\subset\mathbb{R}^{d}$. Without loss of generality, by a scaling argument, we can set $n=1$ outside $\Omega$, so that
$n(\mathbf{x})=\begin{cases}n_{\textup{i}}\in\mathbb{R}_{>0}\quad&\textup{for }\mathbf{x}\in\Omega,\\ 1\quad&\textup{for }\mathbf{x}\notin\Omega,\end{cases}.$ (2)
@@
$-\Delta u_{\textup{inc}}-k^{2}u_{\textup{inc}}=0\quad\textup{in }\mathbb{R}^{d},$ (3)
we can equivalently cast the problem in transmission form:
@@
$u_{\textup{tot}}(\mathbf{x};k)=:\begin{cases}u_{\textup{i}}(\mathbf{x};k)\quad&\textup{for }\mathbf{x}\in\Omega,\\ u_{\textup{s}}(\mathbf{x};k)+u_{\textup{inc}}(\mathbf{x};k)\quad&\textup{for }\mathbf{x}\notin\Omega,\end{cases}$
@@ 2.1 Second-kind boundary-integral equations
Closely following [12, Section 2] and [17, Section 1.3], we rewrite the transmission problem Eq. 4 as a boundary-integral equation. We leverage representations of $u_{\textup{i}}$ and $u_{\textup{s}}$ in terms of layer potentials, via the free-space Green’s function for the Helmholtz operator $\Delta+\kappa^{2}\mathcal{I}$, namely,
$G(\mathbf{x};\kappa):=\begin{cases}\frac{\textup{i}}{4}H_{0}(\kappa\left\|\mathbf{x}\right\|)\quad&\textup{if }d=2,\\ \frac{1}{4\pi\left\|\mathbf{x}\right\|}\exp(\textup{i}\kappa\left\|\mathbf{x}\right\|)\quad&\textup{if }d=3,\\ \end{cases}$
with $H_{0}$ the index-0 Hankel function of first kind. Note, in particular, that the Green’s function depends on $\kappa$ in a very nonlinear but holomorphic way. The (unique) solution of Eq. 4 admits a layer potential representation [12, (5)]
$\begin{cases}u_{\textup{i}}=-\mathsf{V}(kn_{\textup{i}})\big{[}\partial_{\nu}u_{\textup{i}}|_{\Gamma}\big{]}-\mathsf{K}(kn_{\textup{i}})\big{[}u_{\textup{i}}|_{\Gamma}\big{]}\quad&\text{in }\Omega,\\ u_{\textup{s}}=\mathsf{V}(k)\big{[}\partial_{\nu}u_{\textup{s}}|_{\Gamma}\big{]}+\mathsf{K}(k)\big{[}u_{\textup{s}}|_{\Gamma}\big{]}&\text{in }\mathbb{R}\setminus\overline{\Omega},\end{cases}$ (5)
with the “single-layer” and “double-layer” integral operators defined in a parameter-dependent way as
@@
$\mathcal{V}(\kappa)\psi&:=\int_{\Gamma}G(\cdot-\mathbf{y};\kappa)\psi(\mathbf{y})\textup{d}S(\mathbf{y}),\\ \mathcal{K}(\kappa)\phi&:=-\int_{\Gamma}\partial_{\nu(\mathbf{y})}G(\cdot-\mathbf{y};\kappa)\phi(\mathbf{y})\textup{d}S(\mathbf{y}),\\ \mathcal{K}(\kappa)^{*}\phi&:=-\int_{\Gamma}\partial_{\nu(\mathbf{x})}G(\cdot-\mathbf{y};\kappa)\phi(\mathbf{y})\textup{d}S(\mathbf{y}),\\ \mathcal{W}(\kappa)\phi&:=-\int_{\Gamma}\partial_{\nu(\cdot)}\partial_{\nu(\mathbf{y})}G(\cdot-\mathbf{y};\kappa)\phi(\mathbf{y})\textup{d}S(\mathbf{y}).\end{aligned}$ (7)
@@
The former operator $\mathcal{A}(k)$ is bounded on $[L^{2}(\Gamma)]^{2}$ [12, Lemma 1]. Other properties of this formulation of the scattering problem, including its well-posedness, are discussed in Section 3. Another key feature of Eq. 8 is the nonlinear but holomorphic dependence of the two above operators on the wavenumber $k$.
@@
As such, one may sometimes simplify Eq. 8 by replacing $\mathcal{B}(k)$ with the identity operator.
@@ 2.2 Galerkin boundary-element discretization
We consider the boundary-integral equation Eq. 8 in $[L^{2}(\Gamma)]^{2}$ and use the $L^{2}(\Gamma)$ inner product to cast it into a variational form, which is the starting point for Galerkin discretization. Next, we pick a finite-dimensional subspace $\mathcal{X}_{h}$ of $L^{2}(\Gamma)$, with $N_{h}=\dim\mathcal{X}_{h}$, and use $\mathcal{X}_{h}^{2}$ as trial and test spaces. Then we choose a basis $\{\psi_{i}\}_{i=1}^{N_{h}}$ of $\mathcal{X}_{h}$, which induces a basis $\{[\psi_{i},0]^{\top}\}_{i=1}^{N_{h}}\cup\{[0,\psi_{i}]^{\top}\}_{i=1}^{N_{h}}$ of $\mathcal{X}_{h}^{2}$. We use this latter basis to convert Eq. 8 into a $(2N_{h})\times(2N_{h})$ linear system of equations
@@
where the vector $\mathbf{g}$ is defined as the coefficients (with respect to the chosen basis) of the $[L^{2}(\Gamma)]^{2}$-orthogonal projection of $[g_{D},g_{N}]^{\top}$ onto $\mathcal{X}_{h}^{2}$. Moreover, Eq. 9 features the matrix $\textup{B}(k)$, defined entry-wise as
$\textup{B}(k)_{ij}:=\begin{cases}\Big{(}\big{(}{\textstyle\frac{1}{2}}\mathcal{I}-\mathcal{K}(k)\big{)}\psi_{j},\psi_{i}\Big{)}\quad&\text{if }i\leq N_{h}\text{ and }j\leq N_{h},\\ \Big{(}\mathcal{V}(k)\psi_{j-N_{h}},\psi_{i}\Big{)}\quad&\text{if }i\leq N_{h}\text{ and }j>N_{h},\\ \Big{(}\mathcal{W}(k)\psi_{j},\psi_{i-N_{h}}\Big{)}\quad&\text{if }i>N_{h}\text{ and }j\leq N_{h},\\ \Big{(}\big{(}{\textstyle\frac{1}{2}}\mathcal{I}+\mathcal{K}(k)^{*}\big{)}\psi_{j-N_{h}},\psi_{i-N_{h}}\Big{)}\quad&\text{if }i>N_{h}\text{ and }j>N_{h},\\ \end{cases}$
with $(u,v)$ denoting the $L^{2}(\Gamma)$-inner product, as well as $\textup{A}(k)=\textup{M}+\textup{B}(k)-\textup{B}(kn_{\textup{i}})$, with M being the $[L^{2}(\Gamma)]^{2}$-mass matrix, defined entry-wise as
$\textup{M}_{ij}:=\begin{cases}\big{(}\psi_{j},\psi_{i}\big{)}\quad&\text{if }i\leq N_{h}\text{ and }j\leq N_{h},\\ \big{(}\psi_{j-N_{h}},\psi_{i-N_{h}}\big{)}\quad&\text{if }i>N_{h}\text{ and }j>N_{h},\\ 0\quad&\text{otherwise}.\\ \end{cases}$ (10)
@@ 3 Scattering spectrum and field amplification
The well-posedness of the boundary-integral formulation Eq. 8 and of its BE discretization Eq. 9 may be determined by a Fredholm-alternative argument [28, Section 2.1.4], based on the observation that, for all $k\in\mathbb{R}$, $\mathcal{A}(k)$ is a compact perturbation of the identity operator in $[L^{2}(\Gamma)]^{2}$. This makes it an operator of index zero, see, e.g., [30, 28, 12]. Thanks to the injectivity of $\mathcal{A}(k)$ for any nonzero $k$, one can conclude that Eq. 8 is well-posed for all nonzero wavenumbers $k$. From [28, Section 4.2.3] we learn that, provided that $\mathcal{X}_{h}$ possesses sufficiently good $L^{2}(\Gamma)$-approximation properties, the discretized boundary-integral equation Eq. 9 is also well-posed.
@@
obtained by setting $g_{D}=g_{N}=0$, has some relevance on its own. Indeed, it can be interpreted as a nonlinear eigenvalue problem, with $k$ and $[u_{\textup{i}},\partial_{\nu}u_{\textup{i}}]^{\top}$ playing the parts of eigenvalue and eigenfunction, respectively. This problem has been extensively studied, and many of its theoretical properties have been explored before. For instance, eigenvalues must have negative imaginary parts [23, Appendix A], with the exception of 00, which admits the constant eigenfunction $[u_{\textup{i}},\partial_{\nu}u_{\textup{i}}]^{\top}\equiv[1,0]^{\top}$. Moreover, as we show in the next section, the set of eigenvalues is discrete and has no finite accumulation points.

### v2

@@ 1 Introduction
This work is concerned with the study of resonating behavior arising in wave-scattering problems, which, in frequency domain, are modeled through the Helmholtz equation. It is well known from Fredholm theory that, as long as Sommerfeld radiation conditions are enforced, under mild assumptions on the media scattering problems admit a unique solution at any nonzero real frequency. However, in practical terms the scattering problem may no longer be well-posed as a consequence of what we informally call “almost resonances”, namely, frequencies where the inf-sup constant of the problem is dangerously close to 0.
@@
Almost-resonance problems are usually more severe at large frequencies, since, in the large-frequency regime, the complex resonances of the scattering problem get arbitrarily close to the real axis. A good understanding of these phenomena is paramount in many fields. For instance, resonance is sometimes exploited as a way to increase the power of a signal. On the other hand, one may want to avoid resonance to make a physical system more stable or well-behaved. This motivates our interest in this problem.
@@
Incidentally, we also mention that almost resonances have important consequences on the numerical conditioning of the scattering problem after discretization. Thus, spotting almost resonances may be useful for understanding the behavior of numerical methods, e.g., in terms of convergence of iterative linear-system solvers.
@@
At almost-resonating frequencies, the scattered wave may be rather large (in some norm) even for small incident waves. This issue may be equivalently stated in terms of a large operator norm (in suitable spaces) of the solution operator of the scattering problem, which, at fixed frequency, maps an incident wave to a scattered wave. In this work, we use this norm, under the name of “field amplification”, as a quantitative measure of almost-resonating behavior.
@@
For a given scattering problem, our ultimate task is to evaluate the field amplification over a prescribed range of frequency values. A key observation is that the cost of evaluating field amplification over a whole frequency range may be rather high, especially if we use a high resolution both in space and frequency. Indeed, one has to assemble the discrete solution operator and then determine its norm (e.g., by SVD) at each frequency belonging to a sampling grid that covers the target frequency range.
@@
To lessen this burden, we rely on surrogate modeling, which suggests to replace the expensive high-fidelity scattering model with a more economical low(er)-fidelity surrogate. The construction of such surrogate (the so-called “training” phase) is generally not cheap, since it must involve several evaluations of the high-fidelity model. However, it must be carried out only once, independently of the desired number of frequency-domain samples. Ignoring further training costs for simplicity, the complexity of a surrogate-based method for approximating scattering problems is often $\mathcal{O}\left(c_{\textup{high-fidelity}}n+c_{\textup{surrogate}}N\right)$, with $n$ denoting the numbers of training samples and $N$ denoting the number of target frequency evaluations. Compared to the “brute force” approach where the high-fidelity model is evaluated sequentially, computational time can be saved as long as the sampling is parsimonious ($n\ll N$) and the surrogate is efficient ($c_{\textup{surrogate}}\ll c_{\textup{high-fidelity}}$).
@@
In recent years, surrogate modeling has often been applied with the objective of reducing the simulation costs linked to scattering problems [20, 32, 37, 6, 27, 16, 12, 7]. In order to build an effective surrogate, a careful balance usually has to be struck between a priori knowledge of the high-fidelity problem’s structure and a posteriori information obtained from data. Both a priori and a posteriori information must be leveraged in the quest for an efficient method, i.e., keeping $n$ and $c_{\textup{surrogate}}$ low while achieving an acceptable approximation accuracy. The a priori knowledge on the problem is crucial for this, since it is what enables the selection of the “approximation class” within which the surrogate is then sought through a data-driven strategy. In the context of surrogate modeling, the approximation class is crucial for the effectiveness of a numerical algorithm.
@@
However, the effectiveness of the algorithm in [12] is ultimately limited by the fact that the chosen approximation class only partially exploits the structure of the high-fidelity problem. Indeed, the method relies only on the piecewise-smooth nature of the field amplification but not on the meromorphy of the underlying scattering problem, cf. Section 3. In this work, we try to remedy this, by developing methods that better profit from both features. To this aim, we appeal to some key results from operator theory to draw a key connection between our approximation target, namely, field amplification, and a certain nonlinear eigenvalue problem induced by the scattering problem’s differential operator.
A key aspect of our work is that, in building a surrogate for the field amplification, we assume that we can sample at only real wavenumbers. This sets our work apart from, e.g., [3, 6]. Fundamentally, we do this because field amplification is a real quantity, related to real wavenumbers, so it should be possible to approximate it by sampling real wavenumbers only. Moreover, and from a more practical viewpoint, non-real wavenumbers are sometimes out of reach due to limitations in the numerical software that is used to discretize the scattering problem in the first place, and ultimately to compute field amplification. Indeed, adding an imaginary component to the wavenumber is arguably nonphysical, so that, by design, it may not be allowed by, e.g., finite- or boundary-elements libraries.
@@ 2 Boundary-integral equations
The unknown is the total wave $u_{\textup{tot}}:\mathbb{R}^{d}\to\mathbb{C}$, $\textup{i}=\sqrt{-1}$ is the imaginary unit, and $u_{\textup{inc}}:\mathbb{R}^{d}\to\mathbb{C}$ is a given incident wave. Moreover, $n:\mathbb{R}^{d}\to\mathbb{R}_{>0}$ is the index of refraction, which is assumed to attain positive constant values $n_{\textup{i}}$ and $n_{\infty}$ inside and outside, respectively, of a suitable bounded, Lipschitz domain $\Omega\subset\mathbb{R}^{d}$. Without loss of generality, by a scaling argument, we can set $n_{\infty}:=1$ outside $\Omega$, so that
[…]
@@
$\Delta u_{\textup{inc}}-k^{2}u_{\textup{inc}}=0\quad\textup{in }\mathbb{R}^{d},$ (3)
we can equivalently cast the problem in transmission form, denoting by $u_{\textup{i}}$ the total field within $\Omega$ and by $u_{\textup{s}}$ the scattered field outside $\Omega$:
@@
$u_{\textup{tot}}(\mathbf{x};k)=:\begin{cases}u_{\textup{i}}(\mathbf{x};k)\quad&\textup{for }\mathbf{x}\in\Omega,\\ u_{\textup{s}}(\mathbf{x};k)+u_{\textup{inc}}(\mathbf{x};k)\quad&\textup{for }\mathbf{x}\notin\Omega.\end{cases}$
@@ 2.1 Second-kind boundary-integral equations
Closely following [12, Section 2] and [17, Section 1.3], we rewrite the transmission problem Eq. 4 as a boundary-integral equation. We leverage representations of $u_{\textup{i}}$ and $u_{\textup{s}}$ in terms of layer potentials, via the free-space Green’s function for the Helmholtz operator $\Delta+\kappa^{2}$, namely,
[…]
with $H_{0}$ the index-0 Hankel function of first kind. Note, in particular, that the Green’s function depends on $\kappa$ in a very nonlinear but holomorphic way. The (unique) solution of Eq. 4 admits a layer potential representation [12, Eq. (5)]
$\begin{cases}u_{\textup{i}}=-\mathsf{V}(kn_{\textup{i}})(\partial_{\nu}u_{\textup{i}}|_{\Gamma})-\mathsf{K}(kn_{\textup{i}})(u_{\textup{i}}|_{\Gamma})\quad&\text{in }\Omega,\\ u_{\textup{s}}=\mathsf{V}(k)(\partial_{\nu}u_{\textup{s}}|_{\Gamma})+\mathsf{K}(k)(u_{\textup{s}}|_{\Gamma})&\text{in }\mathbb{R}\setminus\overline{\Omega},\end{cases}$ (5)
with the “single-layer” and “double-layer” linear integral operators defined in a parameter-dependent way as
@@
$\begin{aligned} \mathcal{V}(\kappa)\psi&:=\int_{\Gamma}G(\cdot-\mathbf{y};\kappa)\psi(\mathbf{y})\textup{d}S(\mathbf{y}),\\ \mathcal{K}(\kappa)\phi&:=-\int_{\Gamma}\partial_{\nu(\mathbf{y})}G(\cdot-\mathbf{y};\kappa)\phi(\mathbf{y})\textup{d}S(\mathbf{y}),\\ \mathcal{K}(\kappa)^{*}\phi&:=-\int_{\Gamma}\partial_{\nu(\mathbf{x})}G(\cdot-\mathbf{y};\kappa)\phi(\mathbf{y})\textup{d}S(\mathbf{y}),\\ \mathcal{W}(\kappa)\phi&:=-\int_{\Gamma}\partial_{\nu(\cdot)}\partial_{\nu(\mathbf{y})}G(\cdot-\mathbf{y};\kappa)\phi(\mathbf{y})\textup{d}S(\mathbf{y}),\end{aligned}$ (7)
@@
The former operator $\mathcal{A}(k)$ is bounded on $L^{2}(\Gamma)]^{2} [$12, Lemma 1]. Other properties of this formulation of the scattering problem, including its well-posedness, are discussed in Section 3. Another key feature of Eq. 8 is the nonlinear but holomorphic dependence of the two operators $\mathcal{A}(()k)$ and $\mathcal{B}(()k)$ on the wavenumber $k$.
@@
Thus, one may sometimes simplify Eq. 8 by replacing $\mathcal{B}(k)$ with the identity operator.
@@ 2.2 Galerkin boundary-element discretization
We consider the boundary-integral equation Eq. 8 in $L^{2}(\Gamma)]^{2}$ and use the $L^{2}(\Gamma)$ inner product to cast it into a variational form, which is the starting point for Galerkin discretization. Next, we pick a finite-dimensional subspace $\mathcal{X}_{h}$ of $L^{2}(\Gamma)$, with $N_{h}=\dim\mathcal{X}_{h}$, and use $\mathcal{X}_{h}^{2}$ as trial and test spaces. Then we choose a basis $\{\psi_{i}\}_{i=1}^{N_{h}}$ of $\mathcal{X}_{h}$, which induces a basis $\{[\psi_{i},0]^{\top}\}_{i=1}^{N_{h}}\cup\{[0,\psi_{i}]^{\top}\}_{i=1}^{N_{h}}$ of $\mathcal{X}_{h}^{2}$. We use this latter basis to convert Eq. 8 into a $(2N_{h})\times(2N_{h})$ linear system of equations
@@
where the vector $\mathbf{g}$ is defined as the coefficients (with respect to the chosen basis) of the $L^{2}(\Gamma)]^{2}-$orthogonal projection of $g_{D},g_{N}]^{\top}$ onto $\mathcal{X}_{h}^{2}$. Moreover, Eq. 9 features the matrix $\textup{B}(k)$, defined entry-wise as
[…]
with $(u,v)$ denoting the $L^{2}(\Gamma)$-inner product, as well as $\textup{A}(k)=\textup{M}+\textup{B}(k)-\textup{B}(kn_{\textup{i}})$, with M being the $L^{2}(\Gamma)]^{2}-$mass matrix, defined entry-wise as
[…]
@@ 3 Scattering spectrum and field amplification
The well-posedness of the boundary-integral formulation Eq. 8 and of its BE discretization Eq. 9 may be determined by a Fredholm-alternative argument [28, Section 2.1.4], based on the observation that, for all $k\in\mathbb{R}$, $\mathcal{A}(k)$ is a compact perturbation of the identity operator in $L^{2}(\Gamma)]^{2}$. This makes it an operator of index zero, see, e.g., [30, 28, 12]. Thanks to the injectivity of $\mathcal{A}(k)$ for any nonzero $k$ [17, Lemma 1.7], one can conclude that Eq. 8 is well-posed for all nonzero wavenumbers $k$. From [28, Section 4.2.3] we learn that, provided that $\mathcal{X}_{h}$ possesses sufficiently good $L^{2}(\Gamma)$-approximation properties, the discretized boundary-integral equation Eq. 9 is also well-posed.
@@
obtained by setting $g_{D}=g_{N}=0$, has some relevance on its own. Indeed, it can be interpreted as a nonlinear eigenvalue problem, with $k$ and $u_{\textup{i}},\partial_{\nu}u_{\textup{i}}]^{\top}$ playing the parts of eigenvalue and eigenfunction, respectively. This problem has been extensively studied, and many of its theoretical properties have been explored before. For instance, eigenvalues must have negative imaginary parts [23, Appendix A], with the exception of 0, which admits the constant eigenfunction $u_{\textup{i}},\partial_{\nu}u_{\textup{i}}]^{\top}\equiv[1,0]^{\top}$. Moreover, as we show in the next section, the set of eigenvalues is discrete and has no finite accumulation points.